OpenStudy (anonymous):

Solve for x: −9−(9x−6)=3(4x+6) A. 1 B. -1 C. 3 D. -3

OpenStudy (anonymous):

9x-6-9=12x+18 9x-12x=18+15 -3x=33 x=33\-3 x=-11 Or did u mean -9(9x-6) ???